A 100-W lightbulb is placed in a cylinder equipped with a moveable piston. The lightbulb is turned on for 0.020 hour, and the assembly expands from an initial volume of 0.85 L to a final volume of 5.83 Lagainst an external pressure of 1.0 atm.

A) Use the wattage of the lightbulb and the time it is on to calculate ΔE in joules (assume that the cylinder and lightbulb assembly is the system and assume two significant figures).

B) Calculate w.

C) Calculate q.
Answer #1

A) power = energy/time

Power of bulb = 100W

Time for which it was turned on= 0.020 hour = 0.020 × 3600 second = 72 seconds

∆E of system = power × time = 100W × 72 seconds = 7200J

B) work done by the system (w) = P∆V

P= pressure = 1.0atm  

∆V = V​​​​​​final - V​​​​​​initial = 5.83L - 0.85L = 4.98L

w = 1.0atm × 4.98L = 4.98L.atm

1 L.atm = 101.325J

w= 4.98L.atm × 101.325J = 504.6J

C) From first law of thermodynamics ∆E = q + w

q = ∆E - w = 7200J - 504.6J = 6695.5J